OpenStudy (anonymous):

I'm asked to determine if the integral converges or diverges for the triangular region represented by the edges (0,0), (1,1) and (1,2). I can see that the function we integrate has a singularity in (0,0) but does that mean that the actual integral diverges? When I calculated the integral disregarding the singularity it converged.

OpenStudy (anonymous):

This is the integral \[\int\limits_{}^{}\int\limits_{T}^{} \frac{ 1 }{ x \sqrt(y) }\]

OpenStudy (anonymous):

T is the triangular region.